OpenStudy (anonymous):

Please solve this equation step by step 1/2lg(3x+1)-1=lg(x-3)-lg(5)

OpenStudy (anonymous):

is lg(3x+1) in denominator or in numerator? could you retype the question

OpenStudy (anonymous):

in numerator

OpenStudy (anonymous):

did you try? i can give you a hint use LogA-Log B= Log (a/b)

OpenStudy (anonymous):

\[\log_{10} (3x+1)=2* [ \log_{10}[(x-3)/5]\]

OpenStudy (anonymous):

3x+1= [x-3]^2 / 5^2

OpenStudy (anonymous):

25*3x+25=x^2-6x+9

OpenStudy (anonymous):

now solve for x

OpenStudy (anonymous):

thanks ;)

OpenStudy (anonymous):

75x+25=x^2-6x+9